A man is now 30yyears older than his daughter. 15 years ago, he twice as old as his daughter. What are the present ages of the man and his daughter?
Answer by ikleyn(52788)

.

            1.  Mental solution

Since it is given that the man is 30 years older than his daughter,

15 years ago he still was 30 years older than the daughter, 

and, at the same time, was twice her age.


It means that 15 years ago the daughter was 30 years old, while the man was 60 years old.


Hence, now they are 45 and 75 years old.    ANSWER

Solved.


            2.  Algebra solution.

Let x be the daughter age; 

then the man's age is  (x+30) years.


Fifteen years ago the daughter was (x-15) years old, while the man was (x+30)-15 = x+15 years old.


The condition says


    2*(x-15) = x + 15,   or


    2x - 30 = x + 15

    2x - x  = 15 + 30

     x      = 45  years for the daughter,

giving the same answer.

Solved   //   in two ways for your better understanding.

----------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.